Solve 6(|n|- 3)- 4|n|+5=11

Answer by Smartypants077(3)

Solve 6(|n|- 3)- 4|n|+5=11
1. Simplify: 6n +-18+-4n+5=11
2. Combine like terms: 2n+-13=11
3. Seperate terms: 2n=-2
4. Calculate: n=-1
-Bot0789
If this bot is not working to it's potential, feel free to contact the admin.